OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 55 mi / h = ____ ft / s a. 8.7 c. 80.7 b. 26.9 d. 806.7

OpenStudy (anonymous):

here is the unity that changes: miles to feet: \(\frac{5280 ft}{1 mile} \) hours to seconds: \(\frac{1hour}{60second} \)

OpenStudy (anonymous):

now what do I do?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

oops... \(\large \frac{55miles}{1hour} \cdot \frac{5280feet}{1miles} \cdot \frac{1 hour}{3600 seconds} \)

OpenStudy (anonymous):

\(\large \frac{55 \cancel{miles}}{1 \cancel{hour}} \cdot \frac{52870feet}{1 \cancel{miles}} \cdot \frac{1 \cancel{hour}}{3600 seconds} \)